OpenStudy (anonymous):

Find the first four terms of the given sequence tn = n/(n+1). I need the answer..

OpenStudy (anonymous):

Plug in \(n=1,2,3,4\). What do you get?

OpenStudy (anonymous):

Example: \(n=1\) gives you \[t_1=\frac{1}{1+1}=\frac{1}{2}\] This is the first term. What are \(t_2,t_3,t_4\) ?